# Affordance-based Robot Manipulation with Flow Matching

### v4

@@ III-B Flow Matching Policy
We build the robot behavioral cloning policy as a generative process of flow matching, which constructs a flow vector that continuously transforms a source probability distribution toward a destination distribution. Flow matching leverages an ordinary differential equation to deterministically mold data distribution, contrasting with Denoising Diffusion Probabilistic Models (DDPM) that is based on a stochastic differential equation through introducing noise.
@@ III-B1 Flow Matching Model
$p_{0}$ represents a simple base density at time $t=0$, $p_{1}$ denotes the target complicated distribution at time $t=1$, $\bm{x}_{0}$ and $\bm{x}_{1}$ are the corresponding samplings. $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ could be described as:
@@
where we define $\bm{x}_{t}$ as the linear interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with respect to time $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, following the Optimal Transport theory [37]. And $v_{\bm{\theta}}$ is a network of the flow model. Thus Equation (1) could be reformatted as
@@ III-B2 Flow Matching for Visuomotor Policy Learning
Inference: For the inference procedure, random waypoints are sampled from the source distribution and then flowed into the target trajectory by estimating the flow from $t=0$ to $t=1$ over steps. We could adopt multiple steps $1/\Delta t$ for inference:
$\bm{x}_{t+\Delta t}=x_{t}+\Delta tf(\bm{x}_{t},t|\bm{o}),\qquad\text{for}\ t\in[0,1]$ (4)
III-B3 Implementation Details
For the network structures of flow matching, we first adopt ResNet [16] for visual embeddings $\bm{o}$. The flow model $f_{\bm{\theta}}$ is represented with U-Net [40]. The flow model predicts vectors $\mathbf{v}_{t}$ conditioned on visual observation embeddings $\bm{o}$ with Feature-wise Linear Modulation (FiLM) [36] as well as the interpolated waypoints $\bm{x}_{t}$. In the subsequent experiments, we will further study multiple model variants, including transformer-based structure, trajectory representation, etc.
III-B4 Comparisons against Diffusion Policy
In this section, we provide some insights and intuitions about flow matching and its comparisons against diffusion policy for clarification. Algorithm 1 and Algorithm 2 respectively shows the pseudocode of training flow matching and diffusion policy with DDPM. We can see several differences between these two methods:
  • By solving a Stochastic Differential Equation (SDE), DDPM generates a clean sample from Gaussian noise. Flow matching regresses onto a target flow vector field that generates a deterministic mapping from source to target data distributions by solving an ordinary differential equation (ODE).
  • DDPM reverses a diffusion process that adds noise to a clean sample until it becomes Gaussian noise. Flow matching drops all the Gaussian assumptions.
  • Flow matching includes a particularly interesting family of probability paths: the vector field that corresponds to the principle of Optimal Transport with linear interpolant. Flow matching paths with linear interpolation are simpler than diffusion paths, forming straight-line trajectories whereas diffusion paths result in curved paths. These properties seem to empirically translate to more stable training, faster generation, and better performance.
@@ III-C Activities of Daily Living Dataset
We construct a real-world dataset with $10$ tasks across Activities of Daily Living. Each task includes $1,000$ sets of RGB-D images, demonstrated robot action trajectories, and labeled ground truth of affordance maps. Thus $10,000$ demonstrations have been collected in total. The data has been manually collected by moving robot end-effectors with kinesthetic teaching. The novelty of our dataset includes: (i) It contains the same scenarios with multiple objects, multi-task affordances, and the demonstrated robot trajectories. (ii) All tasks are related to Activities of Daily Living that involve (simulated) human data.
We label the affordance heatmaps with 2D Gaussian blobs centered on the object pixels of the demonstrated action. The affordance maps model the locations of all relevant object areas that physically interact with robots given each task. For example, ’feeding task’ requires affordance heatmaps centered on the fork handle, food, and human mouth.
Objects are randomly placed within the range of the table (1.5 meters $\times$ 1 meter). The human (manikin) is placed randomly around the table in the camera view. We have around 30 different objects. Our tasks include prompt primitives: ‘sweep the trash’, ‘pass the water to the human’, ‘hang the towel’, ‘put on the hat’, ‘cover the food’, ‘wipe the nose’, ‘wipe the forearm’, ‘feed the human’, ‘comb the hair’, and ‘brush the teeth’. We can also add a pretrained LLM layer (e. g., GPT) in the very front for zero-shot text classification, allowing for linking other language instructions to one of the ten prompt primitives. The camera is positioned with some variations but generally oriented towards the table and objects. Please refer to the supplementary materials for more videos of each task.
@@ IV Experiments
We systematically evaluate the proposed prompt tuning and flow matching methods against baseline studies. We also ablate how design choices would affect their performance.
@@ IV-A1 Baseline Studies
We benchmark our proposed prompt tuning structures against several commonly used finetuning and instruction-aware vision encoding protocols:
  • Full finetuning: fully update the text and vision transformer layers and the decoder.
  • Adapter-based methods: insert MLP layers with residual connections between pretrained frozen transformer layers of vision and language, as customary in the literature [10].
  • Decoder-based methods: These methods treat the pretrained backbone as a feature extractor with fixed weights during tuning, and only the decoder is tuned, as customary in the literature [16].
  • Side-network methods: train a text transformer network on the side and append pretrained vision features and sidetuned text features before being fed into the decoder, as customary in the literature [9].
  • Cross-attention methods: Similarly to the above side-network methods, the difference here is using cross-attention fusing instead of simple prepending. An example of cross-attention fusing vision and language could be found in the literature [25].
For a fair comparison, all the baselines here use self-supervised pretrained MAE weights on ImageNet-21k dataset for the vision transformer model. We randomly split our Real-world Activities of Daily Living (ADLs) dataset with $80$%-$20$% percentage of training and testing. The results reported here are obtained after $1,000$ epochs of training.
@@ IV-A2 Main Results
Params (M) $\downarrow$ Heatmaps (×$\times 10^{-3}$) $\downarrow$ Centers (pixel) $\downarrow$
Baselines Full $153.8$ $0.76$ $1.15$
Decoder $3.9$ $1.51$ 13.48
Adapter $19.2$ $1.17$ 6.22
Cross-attention $43.5$ $1.26$ 8.89
side-network $42.7$ $1.35$ 9.20
Ours PT-shallow $8.0$ $1.42$ 12.04
PT-deep (self-supervised weights) $42.1$ $0.80$ $2.93$
Ablations PT-deep (supervised weights) $42.1$ $1.48$ 10.13
PT-deep (image output) $42.1$ $1.56$ 13.27
@@
Table I presents the results of prompt tuning on our ADLs testing dataset for affordance learning, comparing against baselines. We use two metrics to evaluate our results: (i) L2 error of affordance heatmap estimation, and (ii) L2 distance between the predicted and ground truth of heatmap centers. We fit Gaussian Mixture Models on predicted heatmaps to determine the inferred heatmap centers. The heatmap error is averaged on each map, and the center error is averaged on per center point. Three observations could be made:
@@
  • Prompt tuning against full finetuning The deep structure of prompt tuning outperforms other baselines except for full finetuning. Full finetuning slightly outperforms deep prompt tuning in terms of heatmap estimation error and heatmap center error. However, the distinction of heatmap center errors ($1.78$ pixels) remains subtle, given the full image size of $224\times 224$. This outcome is favorable as it indicates that most heatmap errors are caused by the tails of the Gaussian distribution, instead of the center area where the robot actions actually applied on. We will further ablate the impact of dataset size on these two methods.
@@
  • Generalizability: We also observe that the trained model could be generalized to new objects. For example, the training dataset only includes a manikin. We found out that it generates well on our testing data with real humans. Affordances on objects with similar shapes (e. g., forks and spoons) could also be transferred. Note that as the proposed tuning method is parameter-efficient, it is envisaged that the method could be readily transferred to different tasks with a small amount of task-specific data.
@@
  • What do prompts learn? We show a t-SNE [46] visualization of the embeddings after the last vision transformer layer (before the decoder) in Fig. 4. We can see that the points of the same color (e. g., tasks with the same language prompts) are embedded together, implying that the representations recover the underlying manifold structure of discriminative task information.
@@
IV-A3 Ablations
We further ablate model design choices:
Pretrained Weights: We evaluate using MAE self-supervised pretrained weights and supervised pretrained weights trained on ImageNet-21k dataset for the vision transformer model. The results in Table I show self-supervised pretrained weights perform better. We are aware of other more complicated variants of vision transformers, for example, CLIP vision encoder and its pretrained weights. As our goal is to integrate textual representations into any vision encoder while keeping it frozen, we have chosen the most basic ViT-B-16 transformer backbone and commonly used pretrained weights and achieved competitive results. We have also tried GPT directly to infer affordance, but no constant performance was achieved without carefully designed prompts or finetuning.
Prompt Location: We have seen different conclusions from prior works about whether the vision-language fusion should be integrated at early or late transformer layers. We conduct experiments to insert prompts at various layers. From Fig. 5-right, we can see that inserting prompts to early layers (for example, layer 1-3 from bottom to top) achieves higher loss than inserting to late layers (for example, layer 1-3 from top to bottom). Thus in our case, prompts have greater significance at the late transformer layers. These results are also supported by the nature of vision transformer hierarchy: lower layers mainly capture low-level fundamental visual details, while higher layers focus on high-level concepts that might be vital for downstream tasks.
@@ IV-B1 Baseline Studies
We compare our flow matching policy against: (i) Diffusion Policy [3] with DDPM and DDIM, and (ii) Transformer-based behavior cloning, as customary in RVT [13], RT-X [35].
Some previous research [34, 11] has mathematically proven that DDIM trajectories are equivalent to flow matching trajectories with a different scaling of time. This scale difference may be beneficial for flow matching. In this experiment, we aim to practically evaluate such speculation. Table V shows the hyperparameters used in flow matching and diffusion policy.
Note that we are aware of other competitive robot behavior cloning methods, including energy-based IBC [7], GAIL [18], etc. Since extensive studies have been conducted and showed better performance of the diffusion policy against these methods, we choose the representative transformer and diffusion baselines for evaluation.
@@
  • Real-world Activities of Daily Living (ADLs) dataset: For 2D data, the training uses a RGB image with visual affordances as input, and the output is a trajectory in 2D pixel space. For the counterparts in 3D, the training takes the concatenation of RGB-D images with visual affordances as input, and outputs a trajectory in 3D Cartesian space. We randomly split the dataset with $80$%-$20$% percentage of training and testing. The results reported are obtained after $1,000$ epochs of training. For each baseline, all 10 tasks in our dataset are trained in a single policy in a supervised manner. To ensure fair comparisons, the corresponding hyperparameters across the flow matching policy and baseline methods are selected to remain consistent. Table VI shows the task summary.
@@
Ours Flow Matching (Transformer) 2.151 4.677 18.05

### v5

@@ III-B Flow Matching Policy
We build the robot behavioral cloning policy as a generative process of flow matching, which constructs a flow vector that continuously transforms a source probability distribution toward a destination distribution. Flow matching leverages an ordinary differential equation to deterministically mold data distribution, contrasting with Denoising Diffusion Probabilistic Models (DDPM), which is based on a stochastic differential equation by introducing noise.
@@ III-B1 Flow Matching Model
$p_{0}$ represents a simple base density at time $t=0$, $p_{1}$ denotes the target complicated distribution at time $t=1$, $\bm{x}_{0}$ and $\bm{x}_{1}$ are the corresponding samplings. $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ is described as:
@@
where we define $\bm{x}_{t}$ as the linear interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with respect to time $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, following the linear conditional flow theory [peyre2019computational]. $v_{\bm{\theta}}$ is a network of the flow model. Thus Equation (1) could be reformatted as
@@ III-B2 Flow Matching for Visuomotor Policy Learning
Inference: For the inference procedure, random waypoints are sampled from the source distribution and then flowed into the target trajectory by estimating the flow from $t=0$ to $t=1$ over steps. We could use multiple steps $1/\Delta t$ for inference:
$\bm{x}_{t+\Delta t}=\bm{x}_{t}+\Delta tf(\bm{x}_{t},t|\bm{o}),\qquad\text{for}\ t\in[0,1]$ (4)
@@
III-B3 Implementation Details
For the network structures of flow matching, we first use ResNet [he2016deep] for visual embeddings $\bm{o}$. The flow model $f_{\bm{\theta}}$ is represented with U-Net [ronneberger2015u]. The flow model predicts vectors $\mathbf{v}_{t}$ conditioned on visual observation embeddings $\bm{o}$ with Feature-wise Linear Modulation (FiLM) [perez2018film] as well as the interpolated waypoints $\bm{x}_{t}$. In the subsequent experiments, we will further study multiple model variants, including a transformer-based structure, trajectory representation.
III-B4 Comparisons against Diffusion Policy
In this section, we provide some insights and intuitions about flow matching and its comparisons against diffusion policy for clarification. Algorithm 1 and Algorithm 2 respectively shows the pseudocode of training flow matching and diffusion policy with DDPM. We can see several similarities and differences between these two methods:
  • Sampling: By solving a Stochastic Differential Equation (SDE), DDPM generates a clean sample from Gaussian noise. Flow matching regresses onto a target flow vector field that generates a deterministic mapping from source to target data distributions by solving an ordinary differential equation (ODE), which drops all the Gaussian assumptions.
  • Probability path: Flow matching includes a particularly interesting family of probability paths: the conditional vector field with linear interpolant. Flow matching paths with linear interpolation are simpler than diffusion paths, forming straighter trajectories, whereas diffusion paths result in curved paths. These properties seem to empirically translate to more stable training, faster generation, and better performance.
  • Reparameterization:  [lu2024simplifying, nakkiran2024step] claims that diffusion models and flow matching can be considered equivalent with reparameterization, including rescaling time and reparameterizing the marginal velocity field via the score function. Specifically,  [nakkiran2024step] argues that DDIM is equivalent to flow-matching with the following parameters:
    $\bm{v}_{t}^{[\bm{x}_{1},\bm{x}_{0}]}(\bm{x}_{t})$ $:=\frac{1}{2t}(\bm{x}_{t}-\bm{x}_{0})$ DDIM
    $\bm{v}_{t}^{[\bm{x}_{1},\bm{x}_{0}]}(\bm{x}_{t})$ $:=\frac{1}{2\sqrt{t}}(\bm{x}_{0}-\bm{x}_{1})$ Flow Matching
    can both generate the same trajectory:
    $\bm{x}_{t}$ $=\bm{x}_{0}+(\bm{x}_{1}-\bm{x}_{0})\sqrt{t}$
    That being said, DDIM at time $t$ corresponds to flow matching at time $2\sqrt{t}$; thus, flow matching is “slower” than DDIM when $t$ is small. This may be beneficial for flow matching in practice.
@@ III-C Activities of Daily Living Dataset
We construct a real-world dataset with $10$ tasks across Activities of Daily Living. Each task includes $1,000$ sets of RGB-D images, demonstrated robot action trajectories, and labeled ground truth of affordance maps. Thus $10,000$ demonstrations have been collected in total. The data has been manually collected by moving robot end-effectors with kinesthetic teaching. The novelty of our dataset includes: (i) It contains the same scenarios with multiple objects, multi-task affordances, and demonstrated robot trajectories. (ii) All tasks are related to Activities of Daily Living that involve (simulated) human data.
We label the affordance heatmaps with 2D Gaussian blobs centered on the object pixels of the demonstrated action. The affordance maps model the locations of all relevant object areas that physically interact with robots, given each task. For example, the feeding task requires affordance heatmaps centered on the fork handle, food, and the human mouth.
Objects are randomly placed within the range of the table (1.5 meters $\times$ 1 meter). The human (manikin) is placed randomly around the table in the camera view. We have around 30 different objects. Our tasks include prompt primitives: ‘sweep the trash’, ‘pass the water to the human’, ‘hang the towel’, ‘put on the hat’, ‘cover the food’, ‘wipe the nose’, ‘wipe the forearm’, ‘feed the human’, ‘comb the hair’, and ‘brush the teeth’. We can also optionally add a pretrained LLM layer (e. g., GPT) in the very front for zero-shot text classification, allowing for linking other language instructions to one of the ten prompt primitives. For example, an ambiguous instruction ‘I am hungry’ can be linked to the prompt primitive ‘feed the human’. The camera is positioned with some variations but generally oriented towards the table and objects. Please refer to the supplementary materials for more videos of each task.
@@ IV Experiments
In this section, we systematically evaluate the performance of the proposed prompt tuning and flow matching methods.
In Section IV-A, we first benchmark our proposed prompt-tuning structures against several commonly used parameter-efficient fine-tuning approaches to demonstrate the superior performance of prompt tuning. We then compare our method with state-of-the-art VLM-based affordance learning approaches to highlight its efficiency. Additionally, we perform ablation studies to analyze the impact of different design choices on performance. In Section IV-B and IV-C, we compare our flow-matching policy with state-of-the-art generative learning, imitation learning, and VLA models for robot manipulation through real-world experiments, evaluating performance in terms of generation quality, stability, inference time, and training resource requirements. In Section IV-D, we systematically investigate the performance of flow matching compared to diffusion policy, across simulation and real-world benchmarks.
@@ IV-A1 Baseline Studies
We benchmark our proposed prompt tuning structures against several commonly used parameter-efficient finetuning protocols and SoA VLM-based affordance learning methods:
  • Full fine-tuning: fully update the text and vision transformer layers and the decoder.
  • Adapter-based methods: insert MLP layers with residual connections between pretrained frozen transformer layers of vision and language, as customary in the literature [liu2023tail, sharma2023lossless].
  • Side-network methods: train a language-based network on the side, append pretrained vision features and sidetuned text features before being fed into the decoder, as customary in the literature [ganz2024question]. This also shares similarities with parallel adaptors. Our deep prompt tuning method differs in the sense that it inserts text layers into every vision layer, while parallel network methods introduce additional trainable modules that run in parallel with the main transformer layers instead of modifying existing layers.
  • Decoder-based methods: adopt the pretrained backbone as a feature extractor with fixed weights during tuning, and only train the decoder, as customary in the literature [he2016deep].
  • Cross-attention methods: use cross-attention fusing text and vision instead of simple prepending. An example of cross-attention fusing vision and language can be found in the literature [jiang2022vima].
  • Mixed-design methods: mixes the favorable adaptor designs. We compare against the state-of-the-art MAM adaptor [he2021towards], which is a Mix-And-Match adapter designed based on the practical findings on NLP.
  • VLM-based affordance learning: We also compare against two state-of-the-art affordance learning methods based on fine-tuning large VLMs: AffordanceLLM [qian2024affordancellm] and Robopoint [yuan2024robopoint].
For a fair comparison, all the baselines here use self-supervised pretrained MAE weights on ImageNet-21k dataset for the vision transformer model. We randomly split our Real-world Activities of Daily Living (ADLs) dataset with 8080%-2020% percentage of training and testing. The results reported here are obtained after $1,000$ epochs of training.
@@ IV-A2 Main Results
Params $\downarrow$ Heatmaps ($\times 10^{-3}$) $\downarrow$ Centers (pixel) $\downarrow$
Baselines Full 153.8M $0.76$ $1.15$
Decoder 3.9M 1.51 13.48
Adapter 19.2M 1.17 6.22
Cross-attention 43.5M 1.26 8.89
side-network 42.7M 1.35 9.20
mixed-design 108M 0.85 2.96
AffordanceLLM 7B $0.72$ $1.01$
Robopoint 12B - 3.18
Ours PT-shallow 8.0M $1.42$ 12.04
PT-deep (self-supervised weights) 42.1M $0.80$ $2.93$
Ablations PT-deep (supervised weights) 42.1M 1.48 10.13
PT-deep (image output) 42.1M 1.56 13.27
@@
We use two metrics to evaluate our results: (i) L2 error of affordance heatmap estimation, and (ii) L2 distance between the predicted and ground truth of heatmap centers. We fit Gaussian Mixture Models on predicted heatmaps to determine the inferred heatmap centers. The heatmap error is averaged on each map, and the center error is averaged on per center point. Three observations could be made:
  • General analysis: Table I presents the results of prompt tuning on our ADLs testing dataset for affordance learning, comparing against baselines. The deep structure of prompt tuning outperforms other parameter-efficient baselines. We can also see that deep prompt tuning achieves better performance than RoboPoint but falls short of the performance achieved by AffordanceLLM. We would like to mention that AffordanceLLM and Robopoint are not parameter-efficient models (the focus of our research), as they involve finetuning large models, including LLaVA and Vicuna. Full finetuning such a large model may not function optimally if only a small dataset is available.
@@
  • Prompt tuning against full finetuning: Full finetuning slightly outperforms deep prompt tuning in terms of heatmap estimation error and heatmap center error. However, the distinction of heatmap center errors ($1.78$ pixels) remains subtle, given the full image size of $224\times 224$. This outcome is favorable as it indicates that most heatmap errors are caused by the tails of the Gaussian distribution, instead of the center area where the robot actions are actually applied. We will further ablate the impact of dataset size on these two methods.
@@
  • Generalizability: We also observe that the trained model could be generalized to new objects. For example, the training dataset only includes a manikin. We found out that it generates well on our testing data with real humans. Affordances on objects with similar shapes (e. g., forks and spoons) could also be transferred. Note that as the proposed tuning method is parameter-efficient, it is envisaged that the method could be readily transferred to different tasks with a small amount of task-specific data. Note that the testing experiments involved only the authors as participants and were therefore exempt from the requirement for ethics approval.
@@
  • What do prompts learn? We show a t-SNE [van2008visualizing] visualization of the embeddings after the last vision transformer layer (before the decoder) in Fig. 4. We can see that the points of the same color (e. g., tasks with the same language prompts) are embedded together, implying that the representations recover the underlying manifold structure of discriminative task information.
  • Prompt tuning or adapters? As pointed by the research of Visual Prompt Tuning [jia2022visual], in contrast to comparable studies in NLP, prompt tuning outperforms full fine-tuning and adapter-based methods in the visual domain. The MAM adaptor [he2021towards] mixes the favorable adaptor designs based on the practical findings on NLP and achieves state-of-the-art results, but does not function optimally in the image-text domain.
IV-A3 Ablations
We further ablate model design choices:
Pretrained Weights: We evaluate using MAE self-supervised pretrained weights and supervised pretrained weights trained on ImageNet-21k dataset for the vision transformer model. The results in Table I show that self-supervised pretrained weights perform better. We are aware of other more complicated variants of vision transformers, for example, CLIP vision encoder and its pretrained weights. As our goal is to integrate textual representations into any vision encoder while keeping it frozen, we have chosen the most basic ViT-B-16 transformer backbone and commonly used pretrained weights and achieved competitive results.
Prompt Location: We have seen different conclusions from prior works about whether the vision-language fusion should be integrated at early or late transformer layers. We conduct experiments to insert prompts at various layers. From Fig. 5-right, we can see that inserting prompts to early layers (for example, layer 1-3 from bottom to top) achieves higher loss than inserting to late layers (for example, layer 1-3 from top to bottom). Thus in our case, prompts have greater significance at the late transformer layers. These results are also supported by the nature of the vision transformer hierarchy: lower layers mainly capture low-level fundamental visual details, while higher layers focus on high-level concepts that might be vital for downstream tasks.
@@ IV-B1 Baseline Studies
We compare our flow matching policy against: (i) Diffusion Policy [chi2023diffusion] with DDPM and DDIM, (ii) Transformer-based behavior cloning, as customary in RVT [goyal2023rvt], RT-X [padalkar2023open], and (iii) Score-based generative models [song2020score] in Section IV-D.
Note that we are aware of other competitive robot behavior cloning methods, including energy-based IBC [florence2022implicit], GAIL [ho2016generative], etc. Since extensive studies have been conducted and showed better performance of the diffusion policy against these methods, we choose the representative transformer and diffusion baselines for evaluation. Table V shows the hyperparameters used in flow matching and diffusion policy.
@@
  • ADLs benchmark: For 2D data, the training uses a RGB image with visual affordances as input, and the output is a trajectory in 2D pixel space. For the counterparts in 3D, the training takes the concatenation of RGB-D images with visual affordances as input, and outputs a trajectory in 3D Cartesian space. Since the ADLs task is not closed-loop, as shown in Table VI, we feed the image with all task-related object affordances to the flow matching policy, which directly outputs the entire trajectory. We randomly split the dataset with 8080%-2020% percentage of training and testing. The results are reported on the testing dataset, with each experiment conducted one time, after $1,000$ epochs of training. For each baseline, all 10 tasks in our dataset are trained in a single policy in a supervised manner. To ensure fair comparisons, the corresponding hyperparameters across the flow matching policy and baseline methods are selected to remain consistent. Table VI shows the task summary.
@@
Ours Flow Matching (Transformer, 16-step) 1.061 1.239 140.71
@@
IV-B2 Main Results
Table II presents the results of flow matching policy on our ADLs testing dataset for robot trajectory learning, comparing against baselines. We use two metrics for evaluation: (i) the average error of 2D and 3D trajectory estimation, and (ii) the average inference time, performed with RTX 4090 GPU. The trajectory error is averaged on each point of the trajectory.
Four observations could be made from this result:
  • Generation quality: Flow matching (CNN-based, 16 steps) outperforms diffusion policy and Transformer baselines in terms of 2D and 3D trajectory prediction accuracy. The suboptimal performance of Transformer behavior cloning is expected as it is hindered by the nature of multi-modal action distribution, causing the averaging out across non-convex spaces.
  • Inference time: We have two observations here: i) Table II showcases that flow matching with 16 steps achieves faster inference time compared to diffusion policy with 16 steps. We hypothesize that flow matching with linear pointwise flows generates straighter flows than DDPM and DDIM, and thus causes faster inference. ii) More importantly, Table II also showcases that 1-step flow matching (error: 1.0311.031cm, time: 13.22813.228ms) has achieved comparable performance as 16-step DDIM (error: 2.2662.266cm, time: 98.68098.680ms), but noticeably lowered inference time roughly by 85%. We hypothesize that this is because diffusion models solve a stochastic differential equation with a series of discrete steps to progressively refine the generated sample. Contrarily, flow matching trains continuously normalize flow models, leading to no significant improvements when increasing inference steps. Thus 2-step flow matching has achieved comparable performance as the 16-step diffusion policy, which considerably reduces the inference time for closed-loop robot manipulation. This is in line with the results obtained in the image generation domain. As pointed out by Stable Diffusion 3 [esser2024scaling], flow matching outperforms diffusion policy with fewer inference steps, making it particularly advantageous in scenarios that demand fast inference.